# Bloch-Floquet band gaps for water waves over a periodic bottom

### v1

@@ 1. Introduction
This study concerns the motion of a free surface wave over a variable bottom. There is a large literature devoted to this subject due to its relevance to oceanography in coastal engineering. Formation of long-shore sandbars along gentle beaches has been observed in open ocean coasts or bays and it is important to understand how they affect the propagation of waves [20]. For mathematical purposes, the variable bottom is often assumed to be periodic or described by a stationary random process. The effect of a fast oscillating bottom has been studied in many asymptotic regimes [28, 5, 9, 6] where effective equations are derived using techniques of homogenization and of multiple scales. Here, we restrict ourselves to the linearized water wave problem near the equilibrium with a periodic bottom. In this setting, there is a classical phenomenon known as the Bragg resonance reflection phenomenon, in analogy with the Bragg’s law for X-rays in crystallography. In the water wave setting, it refers to the situation where the bottom has the form $y=h(x)$, where $h(x)=h_{1}$ (constant) if $x\leqslant 0$ and $x\geqslant\ell$ and is a periodic function for $0<x<\ell$. Resonance happens when the wavelength of the incident wave is equal to twice that of the bottom variation. This configuration leads to strong reflected waves. This phenomenon was observed experimentally by Heathershaw [11] and derived formally by Mei [19] and Miles [21]. Based on this analysis, Mei [19] proposed a theory that strong reflection can be induced by sand bars if the Bragg resonance conditions are met, thus protecting the beach from the full impact of the waves. Higher-order Bragg reflections have been observed experimentally and numerically by Guazzelli, Rey and Belzons [10].
@@ 1.1. Setting of the problem
The problem under consideration is wave propagation over a periodic bottom of infinite extension. Porter and Porter [23] draw a comparison between scattering by a finite length periodic bottom and infinite length periodic bottom and show they are closely related, see also Linton [17], Yu-Howard [31], Liu et al [18], the latter reference dealing with the shallow water limit. The two problems have a different character, the finite extension one giving rise to a boundary-value problem while the infinite extension one takes the form of an eigenvalue problem for the Dirichlet-Neumann operator. The goal of this paper is a detailed study of the Dirichlet-Neumann operator and in particular the description of its spectrum for general periodic topography through an analysis of associated elliptic systems and Bloch-Floquet theory.
@@
The goal of this paper is to describe rigorously the spectrum of the Dirichlet-Neumann operator in the form of bands separated by gaps, thus expressing the wave elevation as a sum of Bloch waves. This decomposition provides the range of frequencies for which propagation of waves is forbidden. Spectral bands correspond to zones of stability while gaps are zones of instability. A classical model operator that exhibits this behavior is the Hill operator $H=-\frac{\,\mathrm{d}}{\,\mathrm{d}x^{2}}+V(x)$, where $V(x)$ is a smooth $2\pi$-periodic potential on ${\mathbb{R}}$. The associated spectral problem is
@@ 1.2. Main results
    The spectrum $\sigma(G[\varepsilon b])$ is purely absolutely continuous and is composed of a union of bands. Namely,
@@
    where the $\{\lambda_{p}^{\varepsilon}(\theta)\}_{p=0}^{\infty}$ are the eigenvalues of $G_{\theta}[\varepsilon b]$, labeled in increasing order, repeated with their order of multiplicity, and the bands are images of the Lipschitz functions $\theta\mapsto\lambda_{p}^{\varepsilon}(\theta)$ on the interval $(-\frac{1}{2},\frac{1}{2}]$. Moreover, $\lambda_{0}^{\varepsilon}(0)=0$.
@@ 3.2. Analyticity of the resolvent of $G_{\theta}[\varepsilon b]$ in $\varepsilon$ and $\theta$
[…]
@@ 3.3. Perturbation theory of self-adjoint operators
    If $A$ has a finite number of eigenvalues in the interval $[a_{1},a_{2}]$ and the sum of their multiplicity is $k$, then this is also true for $A+\varepsilon B_{\varepsilon}$ for all $|\varepsilon|<\varepsilon_{A}$.
@@
$\lambda_{0}^{0}(\theta)<\lambda_{0}^{0}(\tfrac{1}{2})-2d_{p,2}\varepsilon,\quad$ for $\text{for }\theta\in[0,\tfrac{1}{2}-d_{p,1}\varepsilon]$ (3.13)
$\lambda_{2k}^{0}(0)+2d_{p,2}\varepsilon<\lambda_{2k}^{0}(\theta)<\lambda_{2k}^{0}(\tfrac{1}{2})-2d_{p,2}\varepsilon,\quad$ for $\text{for }0<k\leqslant p$
@@
$\lambda_{2k}^{0}(\tfrac{1}{2})+2d_{p,2}\varepsilon<\lambda_{2k+1}^{0}(\theta)<\lambda_{2k+2}^{0}(0)-2d_{p,2}\varepsilon,\quad$ for $\text{for }0\leqslant k\leqslant p$
@@
We now study the nature of the spectrum of $G[\varepsilon b]$. For flat bottom, applying Theorem XIII.86 in [25], we know that the spectrum is purely absolutely continuous and that $\sigma(G[0])=[0,\infty)$ (see below the proof of $\sigma(G[0])=\cup_{k}\lambda_{k}^{0}([-\frac{1}{2},\frac{1}{2}])$), hence $\sigma((1+G[0])^{-1})=[0,1]$. Using the analyticity of the resolvent with respect to $\varepsilon$ (see [15, Appendix A]), we apply Theorem 3.7 for $a_{1}=-\frac{1}{2}$, $a_{2}=\frac{3}{2}$, $A=(1+G[0])^{-1}$ and $\varepsilon B_{\varepsilon}=(1+G[\varepsilon b])^{-1}-(1+G[0])^{-1}$: since $B_{\varepsilon}(A+i)^{-1}$ is bounded uniformly with respect to $\varepsilon$ in $\mathcal{L}(H^{1}({\mathbb{R}}),H^{1}({\mathbb{R}}))$ and $A$ has no eigenvalue in $[a_{1},a_{2}]$, point (b) states that $(1+G[\varepsilon b])^{-1}$ has no eigenvalue in $[a_{1},a_{2}]$ for $|\varepsilon|\leqslant\varepsilon_{A}$, where $\varepsilon_{A}$ depends only on $b$. This implies that $G[\varepsilon b]$ has no eigenvalue. The point (e) of Theorem XIII.85 in [25] infers that for any $\lambda\in{\mathbb{R}}$, $\lambda$ is not an eigenvalue of $G[\varepsilon b]$ if and only if
$\mu(\{\theta\ |\ \lambda\text{ is an eigenvalue of }G_{\theta}[\varepsilon b]\})=0.$
In particular, this gives that $\theta\mapsto\tilde{\lambda}_{n}(\theta,\varepsilon)$ cannot be constant for all $|\varepsilon|<\varepsilon_{A}$. We therefore have all assumptions of Theorem XIII.86 in [25], and we conclude that $G[\varepsilon b]$ has purely absolutely continuous spectrum.
@@ Acknowledgements
This material is based upon work supported by the Swedish Research Council under grant no. 2021-06594 while the authors were in residence at the Institut Mittag-Leffler in Djursholm, Sweden, during the program “Order and Randomness in Partial Differential Equations” in the fall 2023. CL is partially supported by the BOURGEONS project grant ANR-23-CE40-0014-01 of the French National Research Agency (ANR). MM acknowledges financial support from the European Union (ERC, PASTIS, Grant Agreement ${}^{\circ}$101075879). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. CS is partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) under grant no.2018-04536. The authors thank the IDEX Université Grenoble Alpes for funding MM mobility grant.

### v2

@@ 1. Introduction
This study concerns the motion of a free surface wave over a variable bottom. There is a large literature devoted to this subject due to its relevance to oceanography in coastal engineering. Formation of long-shore sandbars along gentle beaches has been observed in open ocean coasts or bays and it is important to understand how they affect the propagation of waves [20]. For mathematical purposes, the variable bottom is often assumed to be periodic or described by a stationary random process. The effect of a fast oscillating bottom has been studied in many asymptotic regimes [28, 5, 9, 6] where effective equations are derived using techniques of homogenization and of multiple scales. Here, we restrict ourselves to the linearized water wave problem near the equilibrium with a periodic bottom. In this setting, there is a classical phenomenon known as the Bragg resonance reflection phenomenon, in analogy with the Bragg’s law for X-rays in crystallography. In the water wave setting, it refers to the situation where the bottom has the form $y=h(x)$, where $h(x)=h_{1}$ (constant) if $x\leqslant 0$ and $x\geqslant\ell$ and is a periodic function of period $d$ for $0<x<\ell$. Resonance happens when the wavelength $\lambda$ of the incident wave is equal to twice that of the bottom variation with higher-order resonances for $\lambda=2d/n,n=2,3$… This configuration leads to strong reflected waves. This phenomenon was observed experimentally by Heathershaw [11] and derived formally by Mei [19] and Miles [21]. Based on this analysis, Mei [19] proposed a theory that strong reflection can be induced by sand bars if the Bragg resonance conditions are met, thus protecting the beach from the full impact of the waves. Higher-order Bragg reflections have been observed experimentally and numerically by Guazzelli, Rey and Belzons [10].
The problem under consideration is wave propagation over a periodic bottom of infinite extension. Porter and Porter [23] draw a comparison between scattering by a finite length periodic bottom and infinite length periodic bottom. The two problems have however a different character, the finite extension one giving rise to a boundary-value problem and the phenomenon of Bragg resonances while the infinite extension one takes the form of an eigenvalue problem. They first observed that the scattering properties for a bottom with an arbitrary number of periods may be obtained from those for the single period through a transfer matrix technique. On the other hand, they defined an extended Bloch problem where they not only look for eigenvalues corresponding to propagating waves over the infinite periodic bottom but also eigenvalues corresponding to evanescent waves, which increase or decrease as they cross a single period. They show that the eigenvalues of the extended transfer matrix used in the scattering case approximate those of the extended Bloch problem (see also Yu and Howard [31]). Other authors pointed out the close relation between the problems of wave propagation over a periodic bottom of finite or infinite extension respectively. We mention the work of Linton [17] who observed a band-gap structure in the context of water waves propagating over infinite periodic arrays of submerged horizontal circular cylinders in deep water. He showed that the approximate location of the band gaps can be obtained from the phase of the transmission coefficient for a single cylinder. In a more recent reference, Liu, Liu and Lin [18], considered shallow water waves over infinite arrays of parabolic bars. They explicitly calculated the band gaps and analyzed the influence of the bar height, width and spacing on band gaps. They found that, if a band gap exists for the spectral problem with an infinite periodic array, it gives rise for the associated problem with a finite periodic extension, to a Bragg resonant reflection occuring at the mid-point of the first band gap, providing a more accurate prediction that the classical Bragg’s law.
@@ 1.1. Setting of the problem
The goal of this paper is a detailed study of the Dirichlet-Neumann operator and in particular the description of its spectrum for general periodic topography through an analysis of associated elliptic systems and Bloch-Floquet theory.
@@
The goal of this paper is to describe rigorously the spectrum of the Dirichlet-Neumann operator in the form of bands separated by gaps, thus expressing the wave elevation as a sum of Bloch waves. This decomposition shows that there are band gaps within which Bloch waves cannot exist. A classical model operator that exhibits this behavior is the Hill operator $H=-\frac{\,\mathrm{d}}{\,\mathrm{d}x^{2}}+V(x)$, where $V(x)$ is a smooth $2\pi$-periodic potential on ${\mathbb{R}}$. The associated spectral problem is
@@ 1.2. Main results
    The spectrum $\sigma(G[\varepsilon b])$ is composed of a union of bands. Namely,
@@
    where the $\{\lambda_{p}^{\varepsilon}(\theta)\}_{p=0}^{\infty}$ are the eigenvalues of $G_{\theta}[\varepsilon b]$, labeled in increasing order, repeated with their order of multiplicity, and the bands are images of the Lipschitz non-negative functions $\theta\mapsto\lambda_{p}^{\varepsilon}(\theta)$ on the interval $(-\frac{1}{2},\frac{1}{2}]$. Moreover, $\lambda_{0}^{\varepsilon}(0)=0$.
@@
  3. (iii)
    The lower part of the spectrum of $G[\varepsilon b]$ is purely absolutely continuous [^1] For a precise definition of the absolutely continuous spectrum we refer to [26, Chapter VII.2].. More precisely, for any $M>0$, there exists $\varepsilon_{M}\in(0,\varepsilon_{0}]$ such that for any $\varepsilon\in[0,\varepsilon_{M})$,
    $\sigma(G[\varepsilon b])\cap[0,M]$ $=\sigma_{\rm ac}(G[\varepsilon b])\cap[0,M]$
    $\sigma_{\rm pp}(G[\varepsilon b])\cap[0,M]$ $=\sigma_{\rm sc}(G[\varepsilon b])\cap[0,M]=\emptyset.$
@@ 3.2. Analyticity of the resolvent of $G_{\theta}[\varepsilon b]$ in $\varepsilon$ and $\theta$
$P(\Sigma)=I_{2}+Q(\Sigma)\quad\text{and}\quad Q(\Sigma)=\varepsilon\begin{pmatrix}-b(x)&zb^{\prime}(x)\\ zb^{\prime}(x)&\frac{b(x)+\varepsilon(zb^{\prime}(x))^{2}}{1-\varepsilon b(x)}\end{pmatrix}.$
@@ 3.3. Perturbation theory of self-adjoint operators
    If $\sigma(A)\cap[a_{1},a_{2}]$ is a set composed of a finite number of eigenvalues and the sum of their multiplicity is $k$, then this is also true for $\sigma(A+\varepsilon B_{\varepsilon})\cap[a_{1},a_{2}]$ for all $|\varepsilon|<\varepsilon_{A}$.
@@
$\lambda_{0}^{0}(\theta)<\lambda_{0}^{0}(\tfrac{1}{2})-2d_{p,2}\varepsilon,$ for $\text{for }\theta\in[0,\tfrac{1}{2}-d_{p,1}\varepsilon]$ (3.13)
$\lambda_{2k}^{0}(0)+2d_{p,2}\varepsilon<\lambda_{2k}^{0}(\theta)<\lambda_{2k}^{0}(\tfrac{1}{2})-2d_{p,2}\varepsilon,$ for $\text{for }0<k\leqslant p$
@@
$\lambda_{2k}^{0}(\tfrac{1}{2})+2d_{p,2}\varepsilon<\lambda_{2k+1}^{0}(\theta)<\lambda_{2k+2}^{0}(0)-2d_{p,2}\varepsilon,$ for $\text{for }0\leqslant k\leqslant p$
@@
We now prove (iii). We first fix $N_{1}$ such that for all $\theta\in[-\frac{1}{2},\frac{1}{2}]$, $\lambda_{N_{1}+1}^{0}(\theta)\geqslant M+2$. By Part (ii) of the theorem, there exists $\varepsilon_{1}(b,N_{1})$ such that for all $\varepsilon\leqslant\varepsilon_{1}(b,N_{1})$ and for all $\theta\in[-\frac{1}{2},\frac{1}{2}]$, $\lambda_{N_{1}+1}^{\varepsilon}(\theta)\geqslant M+1$. Therefore for all $n\geqslant N_{1}+1$,
$\lambda_{n}^{\varepsilon}(\theta)\geqslant M+1.$ (3.14)
Furthermore, there also exists $N$ such that for all $n\geqslant N+1$,
$\tilde{\lambda}_{n}^{\varepsilon}(\theta)\geqslant M+\tfrac{1}{2}.$ (3.15)
We need the above estimate on $\tilde{\lambda}_{n}^{\varepsilon}(\theta)$ in order to reproduce the proof Theorem XIII.86 of [25] and obtain the purely absolutely continuous character of the lower part of the spectrum of $\sigma(G[\varepsilon b])$. To prove (3.15), we proceed by contradiction. Assume that for all $N$, there exist $n\geqslant N+1$ and $\theta_{n}\in[-\frac{1}{2},\frac{1}{2}]$ such that $\tilde{\lambda}_{n}^{\varepsilon}(\theta_{n})<M+\tfrac{1}{2}$, i.e. there is an infinite number of $\tilde{\lambda}_{n}^{\varepsilon}$ passing from $\tilde{\lambda}_{n}^{\varepsilon}(0)=\lambda_{n}^{\varepsilon}(0)>M+1$, to values less that $M+\tfrac{1}{2}$. There is no subsequence such that $(\theta_{n})$ is constant (say equal to $\theta_{0})$, because this would imply that there is an infinite number of eigenvalues at $\theta_{0}$ less than $M+1$. Therefore, we can extract a subsequence $(\theta_{n})$ which is increasing or decreasing, say increasing, where $\tilde{\lambda}_{n}(\theta_{n})<M+\frac{1}{2}$. Denote $\theta_{\infty}\in[-\frac{1}{2},\frac{1}{2}]$ the limit of $(\theta_{n})$. Since there is only a finite number of eigenvalues $(\tilde{\lambda}_{n}^{\varepsilon}(\theta_{\infty}))$ less than $M+1$, an infinite number of $\tilde{\lambda}_{n}^{\varepsilon}$ has to exit $[0,M+\frac{3}{4}]$, i.e. there are $c_{n}\in[\theta_{n},\theta_{\infty})$ such that $M+1\geqslant\tilde{\lambda}_{n}^{\varepsilon}(c_{n})\geqslant M+\frac{3}{4}$. As $\tilde{\lambda}_{n}^{\varepsilon}$ can be defined only from $N_{1}+1$ functions $\lambda_{n}^{\varepsilon}$, there is an index $\varphi(n)\in[0,N_{1}]$ and some values $a_{n},b_{n}$ of the parameter $\theta$ such that, $\theta_{n}\leqslant a_{n}\leqslant b_{n}\leqslant c_{n}$ and $|\lambda_{\varphi(n)}^{\varepsilon}(b_{n})-\lambda_{\varphi(n)}^{\varepsilon}(a_{n})|\geqslant\frac{1}{4(N_{1}+1)}$. As $\varphi(n)\in[0,N_{1}]$, we can extract a sequence such that $\varphi(n)$ is constant (say $N_{\infty}$), which is impossible because $\lambda_{N_{\infty}}^{\varepsilon}$ is Lipschitz. Therefore (3.15) holds.
We define
$H_{n}:=\{h\in L^{2}([-\tfrac{1}{2},\tfrac{1}{2}],\mathbb{T})\;;\;h(\theta)=f(\theta)\tilde{\psi}_{n}^{\varepsilon}(\theta,\varepsilon),f\in L^{2}([-\tfrac{1}{2},\tfrac{1}{2}])\}$ (3.16)
and $G_{n}:H_{n}\longrightarrow H_{n}$ as the multiplication by the function $\tilde{\lambda}_{n}(\cdot,\varepsilon)$.
We use Propositions 3.8, 3.10 and 3.11 to state that there exists $\varepsilon_{2}(N,b)$ such that for all $n\leqslant N$ and for all $\varepsilon\leqslant\varepsilon_{2}(N,b)$ and $\theta\in[-\frac{1}{2},\frac{1}{2}]$, $|\tilde{\lambda}_{n}^{\varepsilon}(\theta)-\tilde{\lambda}_{n}^{0}(\theta)|\leqslant C_{N,b}\varepsilon$. Since $\tilde{\lambda}_{n}^{0}(\theta)=(r+\theta)\tanh(r+\theta)$ for some $r\in\mathbb{N}$, we get that for $\varepsilon$ small enough and for all $n\leqslant N$, $\tilde{\lambda}_{n}^{\varepsilon}(\cdot)$ is not constant. The proof of [25, Theorem XIII.86] then ensures that $G_{n}$ has only absolutely continuous spectrum. Note also that $\sigma(G_{n})=\tilde{\lambda}_{n}^{\varepsilon}([-\frac{1}{2},\frac{1}{2}])$ so by Part (i) of the theorem and (3.15) we have
$\sigma(G[\varepsilon b])\cap[0,M]=\bigcup_{n=0}^{N}\sigma(G_{n})\cap[0,M]$
and therefore (iii).
@@ Acknowledgements
This material is based upon work supported by the Swedish Research Council under grant no. 2021-06594 while the authors were in residence at the Institut Mittag-Leffler in Djursholm, Sweden, during the program “Order and Randomness in Partial Differential Equations” in the fall 2023. CL is partially supported by the BOURGEONS project grant ANR-23-CE40-0014-01 of the French National Research Agency (ANR). MM acknowledges financial support from the European Union (ERC, PASTIS, Grant Agreement n101075879). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. CS is partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) under grant no.2018-04536. The authors thank the IDEX Université Grenoble Alpes for funding MM mobility grant.